OpenStudy (anonymous):

log2 (x)+ logx (4)=3

OpenStudy (anonymous):

i can't understand with this

hero (hero):

Trial and Error seems to be the best solution to this. If you're familiar with the property \(\log_b(b) = 1\) that should help you.

Parth (parthkohli):

\[\log _x 4 = 2 \log _x 2\]Also notice that,\[\log_x 2 = \dfrac{1}{\log_2 x}\]

OpenStudy (anonymous):

assume log log 2 x=t

OpenStudy (anonymous):

t+2/t=3 now solve for t

OpenStudy (anonymous):

then check the solution in the equation

OpenStudy (anonymous):

t^2-3t+2=0

OpenStudy (anonymous):

t=2,1

OpenStudy (anonymous):

log 2 x=2=>x=4 log 2 x=1 => x=2

OpenStudy (kirbykirby):

\[ \large \log_2x+\log_x4=3\\ \large \frac{\log x}{\log2}+\frac{\log 4}{\log x}=3\\ \large \frac{(\log x)^2}{\log 2}+\log 4= 3 \log x\\ \large \frac{(\log x)^2+(\log 2)(\log 4)}{\log 2}= 3 \log x\\ \large (\log x)^2-(3\log 2)(\log x)+(\log 2)(\log 4)=0\\ \large (\log x)^2-(3\log 2)(\log x)+(\log2)(\log2^2)=0\\ \large (\log x)^2-(3\log 2)(\log x)+2(\log 2)^2=0\], since \(\log 2^2= 2 \log 2\) You can think of this a quadratic if you replace \(y=\log x\) \[\large y^2-3\log 2(y)+2(\log 2)^2=0 \] Quadratic formula: \[\large y=\frac{-(-3 \log 2)\pm \sqrt{(-3 \log 2)^2-4(1)(2(\log2)^2)}}{2(1)} \\ \large y=\frac{3 \log 2 \pm \sqrt{9(\log 2)^2 -8(\log 2)^2}}{2}\\ \large y=\frac{3 \log 2 \pm \sqrt{(\log 2)^2}}{2}\\ \large y = \frac{3 \log 2 \pm \log 2}{2}\\ \large y = \frac{4 \log 2}{2}=2 \log 2, \text{ or } \frac{2 \log 2}{2} = \log 2\] Now recall \(y = \log x\) So: \(\large \log x = 2 \log 2 = \log 2^2 = \log 4\) and \(\large \log x = \log 2\) Thus \(\boxed{\large x=2, x=4}\)

Parth (parthkohli):

cody's solution is where I was getting at... but well...